# Dynamics of elastic wires: Preserving Area Without Nonlocality

### v1

@@ Abstract.
We derive a gradient flow of the elastic energy which preserves the enclosed area of evolving planar curves. Contrary to an earlier approach in [40], we give priority to locality, resulting in a sixth order evolution equation. For this new area preserving flow, we prove a global existence result. Additionally, by penalizing the length, we show convergence to an area constrained critical point of the elastic energy.
@@ 1. Introduction
Here, $\nabla_{s}=P^{\perp}\partial_{s}$, where $P^{\perp}$ is the normal projection along $\gamma$ defined by $P^{\perp}\phi=\phi-\langle\phi,\partial_{s}\gamma\rangle\partial_{s}\gamma$ for vector fields $\phi:{\mathbb{S}^{1}}\to\mathbb{R}^{2}$. For $\lambda>0$, solutions to (1.3) have been classified in several works, see for instance [27], [13], or [38, Lemma 5.4]. In particular, the only closed elasticae are multifold coverings of the circle and of the figure eight elastica. For $\lambda=0$, no critical points of (1.2) exist, see e.g. [37, Section 2.2].
@@
In several works, inextensibility of the wire is assumed, i.e. the length of the evolving curve is kept fixed ([25], [7], [14], [44]). In contrast, to the best of our knowledge, there is only one work ([40]) in which the elastic flow under the constraint of fixed enclosed area is studied. The area preserving scenario is physically relevant, for instance, to describe an interface between an inner domain and an outer domain filled with two kinds of incompressible viscous fluids. This relevance is further indicated by the existence of several works exploring area preserving variants for other flows of curves, see for example [18], or the recent articles [47], [24]. In [33, Section 6], the authors explicitly point out the absence of studies of an area preserving elastic flow.
@@
The gradient flow structure ensures that under the flow (1.5), the energy (1.2) decreases (see 3.4). Moreover, the signed enclosed area $\mathcal{A}$ is preserved, i.e.
@@
is constant (see 3.6). Here, $\nu$ denotes the normal vector given by the counter-clockwise rotation of $\partial_{s}\gamma$ by $\pi/2$. Hence, (1.5) is a flow for the elastic energy that preserves the enclosed area without nonlocality. In this article we aim to understand this flow.
@@ 2.1. An appropriate Hilbert space setting
(2.35) $\|\phi\|_{\textup{aux}}=\sup_{Y\in W^{1,2}\setminus\{0\}}\;\frac{1}{\|Y\|_{W^{1,2}}}\int_{\mathbb{S}^{1}}\langle\phi,Y\rangle\,\mathrm{d}s_{\gamma}\leq C\sup_{\begin{subarray}{c}Y\in W^{1,2}\setminus\{0\},\\ Y\perp\gamma\end{subarray}}\;\frac{1}{\|Y\|_{W^{1,2}}}\int_{\mathbb{S}^{1}}\langle\phi,Y\rangle\,\mathrm{d}s_{\gamma}$
@@
for $Y\in W^{1,2}({\mathbb{S}^{1}};\mathbb{R}^{2})$. Moreover, for $Y\in W^{1,2}({\mathbb{S}^{1}};\mathbb{R}^{2})\cap\{Y\perp\gamma\}$, we have
@@
(2.41) $\|\phi\|_{\textup{aux}}\leq C\sup_{\begin{subarray}{c}Y\in W^{1,2},\\ Y\perp\gamma\end{subarray}}\;\frac{1}{\|Y\|_{W^{1,2}}}\int_{\mathbb{S}^{1}}\langle\phi,Y\rangle\,\mathrm{d}s_{\gamma}\leq C\sup_{\begin{subarray}{c}Y\in W^{1,2},Y\perp\gamma,\\ \int_{\mathbb{S}^{1}}\langle Y,\nu_{\gamma}\rangle\,\mathrm{d}s_{\gamma}=0\end{subarray}}\;\frac{1}{\|Y\|_{W^{1,2}}}\int_{\mathbb{S}^{1}}\langle\phi,Y\rangle\,\mathrm{d}s_{\gamma},$
@@ 2.2. The area preserving elastic flow as a gradient flow
Here, we used that
(2.46) $\int_{\mathbb{S}^{1}}\langle C_{\gamma}\nu_{\gamma},\nabla_{s_{\gamma}}^{2}X\rangle\,\mathrm{d}s_{\gamma}=-C_{\gamma}\int_{\mathbb{S}^{1}}\langle\partial_{s_{\gamma}}\nu_{\gamma},\nabla_{s_{\gamma}}X\rangle\,\mathrm{d}s_{\gamma}=0,$
since $\partial_{s_{\gamma}}\nu_{\gamma}$ is tangential. Comparing (2.42) and (2.45), we define the $H_{\gamma}^{-1}$-gradient
@@ 3. Basic properties of the flow
Here, $\nabla_{t}=\partial_{t}^{\perp}=P^{\perp}\partial_{t}$. For a proof of (3.5), (3.6), (3.7), (3.10) and (3.9), see for example [14, Lemma 2.1]. Formula (3.8) follows from (3.7).
As one expects, the flow of 3.2 does not change the rotation index $\omega$ given by
@@
Lemma 3.3.
The rotation index of a sufficiently smooth solution to the flow $\partial_{t}\gamma=V+h\partial_{s}\gamma$, where $V$ is the normal velocity and $h=\langle\partial_{t}\gamma,\partial_{s}\gamma\rangle$, is invariant.
Proof.
Using (3.5), (3.8) and (3.9), we obtain
(3.12) $\frac{\,\mathrm{d}missing}{\,\mathrm{d}t}\int_{\mathbb{S}^{1}}\langle\kappa,\nu\rangle\,\mathrm{d}s=\int_{\mathbb{S}^{1}}\langle\nabla_{s}^{2}V+\langle\kappa,V\rangle\kappa+h\nabla_{s}\kappa,\nu\rangle\,\mathrm{d}s+\int_{\mathbb{S}^{1}}\langle\kappa,\nu\rangle(\partial_{s}h-\langle\kappa,V\rangle)\,\mathrm{d}s.$
With integration by parts and since $\partial_{s}\nu$ is tangential, we see that the right hand side of (3.12) vanishes. ∎
@@
Let $\omega=1$, $\lambda>0$ and $C^{\ast}\approx 146.628$ be as in [36, Theorem 1.1]. Let $\gamma\colon[0,T)\times{\mathbb{S}^{1}}\to\mathbb{R}^{2}$ be a solution to (1.7) with embedded initial datum $\gamma_{0}$ such that
@@
for all $t\in(0,T)$. With [36, Theorem 1.4] we conclude that $\gamma$ describes an embedded curve for all $t\in(0,T)$. ∎
@@ 4. Short time existence
The idea is to to use the reparametrization argument in Section A.1 and to find the solution as a graph over the initial datum. This approach leads to a scalar problem, which we can uniquely solve using A.4.
@@
Step 2: Velocity in terms of $\gamma$. Let $\bar{\gamma}\colon{\mathbb{S}^{1}}\to\mathbb{R}^{2}$ be a sufficiently smooth immersed curve. Continuing the explicit calculation from [44, Proposition A.1], we obtain
@@
for $\tilde{Q}$ a vector valued polynomial. If $\bar{\gamma}$ is time-dependent and $\bar{\gamma}(t,x)=\gamma_{0}(x)+\bar{\varphi}(t,x)\nu_{0}(x)$, then
[…]
@@ 5.2. Bounds on the length
Here, for the second estimate, we used 3.4. If $\lambda>0$, we also have a uniform bound on the length from above by
@@
for all $t\in(0,T)$, where $\varepsilon>0$ can be chosen arbitrarily small. The claim follows with 3.4. ∎
@@ 5.3. Proof of global existence and subconvergence
Putting (5.13), (5.15), (5.16) and (5.17) together, choosing $\varepsilon$ small enough and using 3.4, we get
@@
Step 2: Bound on $\|\partial_{s}^{m}\kappa\|_{C^{0}}$ for $m\in\mathbb{N}_{0}$. By Step 1, 3.4 and 5.4 we have
@@
for $m\in\mathbb{N}_{0}$. As in [6, Theorem 2.2] one shows that
@@
Here, $P(\kappa)$ represents now polynomials of the form $P^{a}_{b}(\kappa)$, $b\in\mathbb{N}$ odd and $7\geq a\in\mathbb{N}$. With this and integration by parts, we see that all terms containing the tangential velocity $\tilde{h}$ in (5.26) cancel. Thus, with the previous steps, $|\frac{\,\mathrm{d}missing}{\,\mathrm{d}t}u(t)|$ is bounded by a constant depending on $\gamma_{0}$ for $t\in[0,\infty)$. Since $u\in L^{1}((0,\infty))$ due to 3.4, this is sufficient to deduce $\lim_{t\to\infty}u(t)=0$. It follows that $\gamma_{\infty}$ solves (1.8) and is therefore stationary. ∎
@@
where $F_{m-1}$ is a polynomial of degree at most $m-1$. With Step 2, it therefore suffices to find a bound on $\|\partial_{x}^{m}|\partial_{x}\gamma|\|_{C^{0}}$ for $m\in\mathbb{N}_{0}$. For this, we proceed by induction as in [10, p. 649]. For $m=0$, we note that $\partial_{t}|\partial_{x}\gamma|=-\langle\kappa,V\rangle|\partial_{x}\gamma|$ by (3.5). Since $\|\langle\kappa,V\rangle\|_{C^{0}}$ is bounded by (5.22) and (5.23) and $|\partial_{x}\gamma_{0}|$ is bounded from below and above, we conclude that
@@ 6.1. Stationary solutions as constrained critical points
for all $Y\in W^{2,2}({\mathbb{S}^{1}};\mathbb{R}^{2})$. For a computation of $\mathcal{E}^{\prime}(\gamma)$, see for example [10, Lemma A.1]. If $\gamma$ is sufficiently regular, we obtain from (6.3) the Euler Lagrange equation
@@ 6.2. A constrained Łojasiewicz–Simon gradient inequality
As in [11] or in [44], we first prove a Łojasiewicz–Simon gradient inequality in $V^{\perp}$, i.e. for normal directions. Since we work in a constrained setting, the inequality will be based on [43]. The following two lemmata cover the assumptions in [43, Cor. 5.2].
@@
For the definition and fundamental properties of analytic maps between Banach spaces, see for example [51]. A detailed proof of 6.2 in the case of open elastic curves can be found in [11, Section 3]. By slightly changing the function spaces (no boundary conditions but working in ${\mathbb{S}^{1}}$), this proof is transferred to the case of closed curves. A proof of 3. in 6.2 is also given in [34]. We remark that for this Fredholm property of the second variation, it is crucial that we consider only variations in normal direction.
@@ 3. With the projection $P^{\perp}=\langle\cdot,\nu_{\hat{\gamma}}\rangle\nu_{\hat{\gamma}}$, we compute
(see for example [11, (B.5)]), which implies that the first summand in (6.17) vanishes. For the second summand, we used
@@
(see for example [11, (B.2)]). We see that $(dG^{\prime})(0)\colon V^{\perp}\to L^{2,\perp}$ is of zeroth order in $Y$. Hence it is compact.
@@
Let $\hat{\gamma}\in C^{\infty}_{\mathrm{Imm}}({\mathbb{S}^{1}};\mathbb{R}^{2})$ be a constrained critical point of $\mathcal{E}_{\lambda}$. Then there exists constants $C>0$, $\vartheta\in(0,\frac{1}{2}]$ and $r>0$, such that
@@
We check that the assumptions in [43, Corollary 5.2] are satisfied for $E_{\lambda}$ and $G$ defined on $U_{\varepsilon}\subset V^{\perp}$. In 6.2 and 6.3 we have seen that $E$ and $G$ are analytic. Assumption (i), i.e. $V^{\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})\hookrightarrow L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})$ densely, is given. Assumptions (ii) and (iii) follow from 6.2 and assumptions (iv) to (vi) from 6.3. Thus, by [43, Corollary 5.2], the set of area preserving normal variations
@@
is locally a submanifold of codimension $1$ near 00. Since $\hat{\gamma}$ is a constrained critical point of $\mathcal{E}_{\lambda}$, 00 is a critical point of $E_{\lambda}$ on $\mathcal{M}$. Thus, [43, Corollary 5.2] further yields the existence of constants $C,r>0$ and $\vartheta\in(0,\tfrac{1}{2}]$ such that for all $Y\in\mathcal{M}$ with $\|Y\|_{W^{4,2}(\,\mathrm{d}x)}\leq r$, it holds that
@@
where $P(Y)\colon L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})\to L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})$ is the orthogonal projection onto the closure of the tangent space $\overline{T_{Y}\mathcal{M}}^{\|\cdot\|_{L^{2}(\,\mathrm{d}x)}}$. With [43, Proposition 3.3] it follows that
@@
By [43, Remark 5.4], the projection operator $P(Y)\colon L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})\to L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})$ is given by

### v2

@@ Abstract.
We derive an $H^{-1}$-gradient flow of the elastic energy which preserves the enclosed area of evolving planar curves. For this new sixth-order evolution equation, we prove a global existence result. Additionally, by penalizing the length, we show convergence to an area constrained critical point of the elastic energy.
@@ 1. Introduction
Here, $\nabla_{s}=P^{\perp}\partial_{s}$, where $P^{\perp}$ is the normal projection along $\gamma$ defined by $P^{\perp}\phi=\phi-\langle\phi,\partial_{s}\gamma\rangle\partial_{s}\gamma$ for vector fields $\phi:{\mathbb{S}^{1}}\to\mathbb{R}^{2}$. For $\lambda>0$, solutions to (1.3) have been classified in several works, see for instance [29], [15], or [40, Lemma 5.4]. In particular, the only closed elasticae are multifold coverings of the circle and of the figure eight elastica. For $\lambda=0$, no critical points of (1.2) exist, see e.g. [39, Section 2.2].
@@
In several works, inextensibility of the wire is assumed, i.e. the length of the evolving curve is kept fixed ([27], [9], [16], [46]). In contrast, to the best of our knowledge, there is only one work ([42]) in which the elastic flow under the constraint of fixed enclosed area is studied. The area preserving scenario is physically relevant, for instance, to describe an interface between an inner domain and an outer domain filled with two kinds of incompressible viscous fluids. This relevance is further indicated by the existence of several works exploring area preserving variants for other flows of curves, see for example [20], or the recent articles [49], [26]. In [35, Section 6], the authors explicitly point out the absence of studies of an area preserving elastic flow.
@@
The gradient flow structure ensures that under the flow (1.5), the energy (1.2) decreases (see 3.3). Moreover, the signed enclosed area $\mathcal{A}$ is preserved, i.e.
@@
is constant (see 3.5). Here, $\nu$ denotes the normal vector given by the counter-clockwise rotation of $\partial_{s}\gamma$ by $\pi/2$. Hence, (1.5) is a flow for the elastic energy that preserves the enclosed area without nonlocality. In this article we aim to understand this flow.
@@ 2.1. An appropriate Hilbert space setting
(2.35) $\|\phi\|_{\textup{aux}}=\sup_{Y\in W^{1,2}\setminus\{0\}}\;\frac{1}{\|Y\|_{W^{1,2}}}\int_{\mathbb{S}^{1}}\langle\phi,Y\rangle\,\mathrm{d}s_{\gamma}\leq C\sup_{\begin{subarray}{c}Y\in W^{1,2}\setminus\{0\},\\ Y=\langle Y,\nu_{\gamma}\rangle\nu_{\gamma}\end{subarray}}\;\frac{1}{\|Y\|_{W^{1,2}}}\int_{\mathbb{S}^{1}}\langle\phi,Y\rangle\,\mathrm{d}s_{\gamma}$
@@
for $Y\in W^{1,2}({\mathbb{S}^{1}};\mathbb{R}^{2})$. Moreover, for $Y\in W^{1,2}({\mathbb{S}^{1}};\mathbb{R}^{2})\cap\{Y=\langle Y,\nu_{\gamma}\rangle\nu_{\gamma}\}$, we have
@@
(2.41) $\|\phi\|_{\textup{aux}}\leq C\!\sup_{\begin{subarray}{c}Y\in W^{1,2}\setminus\{0\},\\ Y=\langle Y,\nu_{\gamma}\rangle\nu_{\gamma}\end{subarray}}\;\frac{1}{\|Y\|_{W^{1,2}}}\!\int_{\mathbb{S}^{1}}\langle\phi,Y\rangle\,\mathrm{d}s_{\gamma}\leq C\!\sup_{\begin{subarray}{c}Y\in W^{1,2}\setminus\{0\},Y=\langle Y,\nu_{\gamma}\rangle\nu_{\gamma},\\ \int_{\mathbb{S}^{1}}\langle Y,\nu_{\gamma}\rangle\,\mathrm{d}s_{\gamma}=0\end{subarray}}\;\frac{1}{\|Y\|_{W^{1,2}}}\!\int_{\mathbb{S}^{1}}\langle\phi,Y\rangle\,\mathrm{d}s_{\gamma},$
@@ 2.2. The area preserving elastic flow as a gradient flow
Here, we used in the third step that
(2.46) $\int_{\mathbb{S}^{1}}\langle C_{\gamma}\nu_{\gamma},\nabla_{s_{\gamma}}^{2}X\rangle\,\mathrm{d}s_{\gamma}=0,$
since $\nabla_{s_{\gamma}}^{2}X=-\phi\in H_{\gamma}$. Comparing (2.42) and (2.45), we define the $H_{\gamma}^{-1}$-gradient
@@ 3. Basic properties of the flow
Here, $\nabla_{t}=\partial_{t}^{\perp}=P^{\perp}\partial_{t}$. For a proof of (3.5), (3.6), (3.7), (3.9) and (3.10), see for example [16, Lemma 2.1]. Formula (3.8) follows from (3.7).
Clearly, the flow of 3.2 does not change the rotation index $\omega$ given by
@@
Let $\omega=1$, $\lambda>0$ and $C^{\ast}\approx 146.628$ be as in [38, Theorem 1.1]. Let $\gamma\colon[0,T)\times{\mathbb{S}^{1}}\to\mathbb{R}^{2}$ be a solution to (1.7) with embedded initial datum $\gamma_{0}$ such that
@@
for all $t\in(0,T)$. With [38, Theorem 1.4] we conclude that $\gamma$ describes an embedded curve for all $t\in(0,T)$. ∎
@@ 4. Short time existence
The idea is to use the reparametrization argument in Section A.1 and to find the solution as a graph over the initial datum. This approach leads to a scalar problem, which we can uniquely solve using A.4.
@@
Step 2: Velocity in terms of $\gamma$. Let $\bar{\gamma}\colon{\mathbb{S}^{1}}\to\mathbb{R}^{2}$ be a sufficiently smooth immersed curve. Continuing the explicit calculation from [46, Proposition A.1], we obtain
@@
for $\tilde{Q}$ a vector valued polynomial. If $\bar{\gamma}$ is time-dependent and $\bar{\gamma}(t,x)=\gamma_{0}(x)+\bar{\varphi}(t,x)\nu_{0}(x)$, then $P^{\perp_{\bar{\gamma}}}\partial_{x}^{6}\bar{\gamma}$ consists of the term $\partial_{x}^{6}\bar{\varphi}P^{\perp_{\bar{\gamma}}}\nu_{0}$ and of terms containing derivatives of $\varphi$ up to order 5.
@@ 5.2. Bounds on the length
Here, for the second estimate, we used 3.3. If $\lambda>0$, we also have a uniform bound on the length from above by
@@
for all $t\in(0,T)$, where $\varepsilon>0$ can be chosen arbitrarily small. The claim follows with 3.3. ∎
@@ 5.3. Proof of global existence and subconvergence
Putting (5.13), (5.15), (5.16) and (5.17) together, choosing $\varepsilon$ small enough and using 3.3, we get
@@
Step 2: Bound on $\|\partial_{s}^{m}\kappa\|_{C^{0}}$ for $m\in\mathbb{N}_{0}$. By Step 1, 3.3 and 5.4 we have
@@
for $m\in\mathbb{N}_{0}$. As in [7, Theorem 2.2] one shows that
@@
Here, $P(\kappa)$ represents now polynomials of the form $P^{a}_{b}(\kappa)$, $b\in\mathbb{N}$ odd and $7\geq a\in\mathbb{N}$. With this and integration by parts, we see that all terms containing the tangential velocity $\tilde{h}$ in (5.26) cancel. Thus, with the previous steps, $|\frac{\,\mathrm{d}missing}{\,\mathrm{d}t}u(t)|$ is bounded by a constant depending on $\gamma_{0}$ for $t\in[0,\infty)$. Since $u\in L^{1}((0,\infty))$ due to 3.3, this is sufficient to deduce $\lim_{t\to\infty}u(t)=0$. It follows that $\gamma_{\infty}$ solves (1.8) and is therefore stationary. ∎
@@
where $F_{m-1}$ is a polynomial of degree at most $m-1$. With Step 2, it therefore suffices to find a bound on $\|\partial_{x}^{m}|\partial_{x}\gamma|\|_{C^{0}}$ for $m\in\mathbb{N}_{0}$. For this, we proceed by induction as in [12, p. 649]. For $m=0$, we note that $\partial_{t}|\partial_{x}\gamma|=-\langle\kappa,V\rangle|\partial_{x}\gamma|$ by (3.5). Since $\|\langle\kappa,V\rangle\|_{C^{0}}$ is bounded by (5.22) and (5.23) and $|\partial_{x}\gamma_{0}|$ is bounded from below and above, we conclude that
@@ 6.1. Stationary solutions as constrained critical points
for all $Y\in W^{2,2}({\mathbb{S}^{1}};\mathbb{R}^{2})$. For a computation of $\mathcal{E}^{\prime}(\gamma)$, see for example [12, Lemma A.1]. If $\gamma$ is sufficiently regular, we obtain from (6.3) the Euler Lagrange equation
@@ 6.2. A constrained Łojasiewicz–Simon gradient inequality
As in [13] or in [46], we first prove a Łojasiewicz–Simon gradient inequality in $V^{\perp}$, i.e. for normal directions. Since we work in a constrained setting, the inequality will be based on [45]. The following two lemmata cover the assumptions in [45, Cor. 5.2].
@@
For the definition and fundamental properties of analytic maps between Banach spaces, see for example [53]. A detailed proof of 6.2 in the case of open elastic curves can be found in [13, Section 3]. By slightly changing the function spaces (no boundary conditions but working in ${\mathbb{S}^{1}}$), this proof is transferred to the case of closed curves. A proof of 3. in 6.2 is also given in [36]. We remark that for this Fredholm property of the second variation, it is crucial that we consider only variations in normal direction.
@@ 3. With the projection $P^{\perp}=\langle\cdot,\nu_{\hat{\gamma}}\rangle\nu_{\hat{\gamma}}$, we compute
(see for example [13, (B.5)]), which implies that the first summand in (6.17) vanishes. For the second summand, we used
@@
(see for example [13, (B.2)]). We see that $(dG^{\prime})(0)\colon V^{\perp}\to L^{2,\perp}$ is of zeroth order in $Y$. Hence it is compact.
@@
Let $\hat{\gamma}\in C^{\infty}_{\mathrm{Imm}}({\mathbb{S}^{1}};\mathbb{R}^{2})$ be a constrained critical point of $\mathcal{E}_{\lambda}$. Then there exist constants $C>0$, $\vartheta\in(0,\frac{1}{2}]$ and $r>0$, such that
@@
We check that the assumptions in [45, Corollary 5.2] are satisfied for $E_{\lambda}$ and $G$ defined on $U_{\varepsilon}\subset V^{\perp}$. In 6.2 and 6.3 we have seen that $E$ and $G$ are analytic. Assumption (i), i.e. $V^{\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})\hookrightarrow L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})$ densely, is given. Assumptions (ii) and (iii) follow from 6.2 and assumptions (iv) to (vi) from 6.3. Thus, by [45, Corollary 5.2], the set of area preserving normal variations
@@
is locally a submanifold of codimension $1$ near 00. Since $\hat{\gamma}$ is a constrained critical point of $\mathcal{E}_{\lambda}$, 00 is a critical point of $E_{\lambda}$ on $\mathcal{M}$. Thus, [45, Corollary 5.2] further yields the existence of constants $C,r>0$ and $\vartheta\in(0,\tfrac{1}{2}]$ such that for all $Y\in\mathcal{M}$ with $\|Y\|_{W^{4,2}(\,\mathrm{d}x)}\leq r$, it holds that
@@
where $P(Y)\colon L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})\to L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})$ is the orthogonal projection onto the closure of the tangent space $\overline{T_{Y}\mathcal{M}}^{\|\cdot\|_{L^{2}(\,\mathrm{d}x)}}$. With [45, Proposition 3.3] it follows that
@@
By [45, Remark 5.4], the projection operator $P(Y)\colon L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})\to L^{2,\perp}({\mathbb{S}^{1}};\mathbb{R}^{2})$ is given by